The Grand Multiplier:
c = 0.6
t = 0.2
m = 0.1
AEo = 20,000.
Determine Y...
1
Expert's answer
2016-05-11T12:36:03-0400
c = 0.6
t = 0.2
m = 0.1
AEo = 20,000.
AE = C = c*Y = AEo + c*(Y - T) = AEo + c*(1 - t)*Y,
0.6Y = 20,000 + 0.6*(1 - 0.2)*Y,
0.12Y = 20,000,
Y = 20,000/0.12 = 166,666.67.